OpenStudy (anonymous):

Brandon needs $480 to buy a TV and stereo system for his room. He received $60 in cash for birthday presents. He plans to save $30 per week from his part-time job. To find how many weeks w it will take to have $480, solve 60 + 30w = 480. Type your answer in the blank to the left. A. 16 weeks B. 13 weeks C. 15 weeks D. 14 weeks

OpenStudy (anonymous):

try all the answers plug in for w

OpenStudy (anonymous):

just take 60, move to the other side than divide by 30 and there you go

OpenStudy (anonymous):

its D 14 weeks

OpenStudy (anonymous):

yeap i get that thanks

OpenStudy (anonymous):

no problem

OpenStudy (texaschic101):

60 + 30w = 480 30w = 480 - 60 30w = 420 w = 14